QUESTION TWO An economy is represented by the following set of equations: Y=C+I C = 80m + 16Y where Y represents aggregate exQUESTION TWO An economy is represented by the following set of equations: Y = C + I C = 80m + 1.6Y where Y represents aggregate expenditure C represents consumption expenditure by households I represents investment expenditure by firms M is millions of Ghana cedis (a) (i) Explain why the model is not an open economy? (2 marks) (ii) Explain investment expenditure (I) as used in the model. (3 marks) (b) Using the consumption function, estimate the autonomous consumption; marginal propensity to consume (MPC) and the multiplier. (6 marks) (c) Given that the value of Investment (I) is GHȼ100m calculate equilibrium (i) Estimate the aggregate expenditure (Y) (3 marks) (ii) Consumption expenditure (C) (3 marks) (iii) Aggregate saving (S) (3 marks)
Answer #1

2. a)i) This model is not an open economy becuase there are no exports and imports. The Net exports component is not included in the model indicating a closed economy.

ii) Investment expenditure in the model is a fixed (autonomous) value in millions of Ghana cedis. This investment spending is not a function of interest rate or output.

b) The consumption function is of the form C = C(bar) + mpc. Y. Given the consumption function is in the same form, we can eastimate that the autonomous consumption (C(bar)) is equal to 80million while mpc is 1.6 (This is impractical- mpc lies between 0 to 1. This number will probably be 0.6). (Remember mpc+mps=1)

The multiplier is given by 1/mps. Given mpc = 1.6, multiplier = 1/1-1.6 = 1/-0.6 = -1.666 (impractical result)

(If my assumption is correct and the mpc is infact 0.6 then multiplier will be 1/0.4 = 2.5)

c)iAggregate expenditure Y = C+I

= 80m + 1.6Y + 100m

= 180m + 1.6Y

at equilibrium, Y= 180m + 1.6Y

or, -0.6Y = 180m

or, Y = (-300m) (impractical result because mpc given is incorrect.)

At equilibrium, Y=AE=-300m.

When mpc = 0.6, Y = 180m + 0.6Y

at equilibrium, 0.4Y = 180m; Y=450m

ii) For the next two sub parts I will consider that mpc=1.6 is correct. So then Y=-300m

C= 80m - 480m = -400m

iii) Y= C+S

S = Y-C = -300m - 400m = -700m

(Impractical results)

Considering mpc = 0.6, Y=450m

C= 80m + 0.6*450m = 350m

S= 450m-350m =100m (correct results as S=I in closed economy)